The attendance at a moving theater on a certain day was 737 persons. If there were 289 more adults than children, how many children were there?
Found 2 solutions by checkley77, Jeannie1012:
Answer by checkley77(12844)

C+C+289=737
2C=737-289
2C=448
C=448/2
C=224 CHILDREN.
224+289=513 ADULTS.
PROOF:
224+513=737
737=737

Answer by Jeannie1012(8)

You are going to use substitution to answer this problem.
First make clear what will represent what. You write a let statement for this.
Step 1:Let a= adults and Let c = children
we know that there are 737 people total so we write the equation
step 2: a + c = 737
we know that there are 289 more adults than children. So adults a minus 289 = children c. so you substitute that into your equation a + (a-289) = 737
step 3: a+a = 2a and you can't add letters and number so it become 2a-289 =737
step 4: 2a -289 = 737 add 289 to both sides to get the a on a side by its self 2a = 1026
step 5: divide both sides by 2 to get the a alone
2a /2 = a and 1026/2 = 513
so a = 513
Step 6: plug it in to the orginal a + c = 737
513 + c = 737
subtract 513 from both sides c = 224
now you have the value of both a and c you can plug them in and check
a + c = 737
513 + 224 =737
so there were 224 children and 513 adults.